# Learning from galactic rotation curves: a neural network approach

### v1

@@ I Introduction
These novel tools and techniques are currently being applied to a wide range of astrophysical and cosmological datasets [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. For instance, Refs. [9, 14] use ANNs to compute likelihood in Bayesian inference to reduce computational time. On the other hand, Ref. [10] carries out non-parametric reconstruction of the Hubble parameter as a function of redshift while Ref. [18] does so for galactic rotational velocity as a function of radius. Ref. [10] finds that if one conducts Bayesian inference using Markov Chain Monte-Carlo (MCMC) on the reconstructed data, the resultant cosmological parameter posteriors are in agreement to those obtained from observed data. Deep learning has also been applied to the reconstruction of CMB B-modes[19], as well as to reconstruct full CMB spectra from partial sky data [16, 17]. Neural networks have also been used to estimate parameters from observational data like $H(z)$ data [15], CMB angular power spectrum [11], CMB birefringence maps [20] as well as Lyman-$\alpha$ spectra [21].
In this work, we use neural networks to extract model parameters from galactic rotation curves from the SPARC catalog [24]. Recall that rotation curves are pivotal in tracing the mass distribution of dark matter and baryons in galaxies, and are an important test of dark matter models [25]. We consider dark matter to comprise of a spin-zero particle with mass $m\sim 10^{-22}\ \text{eV}$ (called Ultra-Light Dark Matter (ULDM)), whose large deBroglie wavelength leads to the formation of a flat density core surrounded by a cold dark matter-like envelope [26] (see also [27, 28, 29]). We thus have the following five free parameters: mass of the dark matter particle $m\ (\text{eV})$, along with galaxy specific parameters such as the scaling parameter $s$, which characterizes the dark matter core, core-to-envelope transition radius $r_{t}\ (\text{kpc})$ and NFW scale radius $r_{s}\ (\text{kpc})$ which characterize the surrounding halo. The effect of Baryons is parameterized by the stellar mass-to-light ratio $\Upsilon_{*}\ (M_{\odot}/L_{\odot})$, which tunes contribution from the stellar disk (see Section II.1 for details).
@@
We explore whether, using simulated rotation curves, one can train a neural network to approximate this function. Typically, $N_{obs}\sim\mathcal{O}(15)$ for the galaxies we consider, while the number of parameters we want to infer is $5$ (or $10$ if uncertainties are also inferred). If the neural network has $2$ hidden layers with $200$ neurons each, it will have $\sim(40-50)\times 10^{3}$ internal adjustable parameters (called weights and biases). To fix these internal parameters, we need to train the neural network. For training, we use simulated rotation curves whose parameter values are already known. We generate training data for $7$ dwarf galaxies from the SPARC catalog [24] and train a different neural network for each galaxy. The size of the training data, i.e., the number of known pairs of rotation curves and parameters in our work is $\sim 10^{5}$. The details of neural networks used and their architectures are discussed in Sections II.2.1 and II.5 respectively.
To test our trained neural networks, we use the observed rotation curves for the $7$ galaxies as input and infer parameter values in Section III.1. Then, in Section III.2, we explore the effect of noise in the training data on performance of the neural network during parameter inference, and find that the including noise improves point-estimates of parameters when confronted with observed rotation curves, i.e., the rotation curve obtained from these parameter values agree well with the observed rotation curves. In Section IV, we also utilize two different ways of obtaining uncertainties in the model parameters: during inference (as carried out in [11]), or during training (following the work in [15]). Finally, we compare the parameter point-estimates and uncertainties obtained using our approach to those obtained using MCMC in Section V. We conclude in Section VI.
II Rotation curves and artificial neural networks
II.1 Model and data
Galactic rotation curves, i.e. orbital velocity of stars and gas as a function of distance from the centers of galaxies are an important probe of the matter (visible and dark) distribution in said galaxies.
@@ II.1 Model and data
$V_{obs}=\sqrt{V_{DM}^{2}+V_{g}|V_{g}|+\Upsilon_{d}V_{d}|V_{d}|+\Upsilon_{b}V_{b}|V_{b}|}\ ,$ (1)
@@
In this work, we consider dark matter to comprise of ultralight spin-zero scalars, with $m\sim 10^{-22}\ \text{eV}$. Due to the large deBroglie wavelength $\mathcal{O}(1\ \text{kpc})$, simulations suggest that dark matter halos in the ULDM paradigm have a core-halo structure where, the inner regions of the halo are described by flat density cores [26]. These cores are stationary state solutions of the Schrödinger-Poisson system of equations. In the outer regions, beyond a transition radius, ULDM behaves like CDM and the corresponding density profile can be described by the well known Navarro-Frenk-White (NFW) profile [35]. Hence, for a galactic halo, the total dark matter density profile can be written as
$\rho_{DM}(r)=\rho_{ULDM}\Theta(r_{t}-r)+\rho_{NFW}\Theta(r-r_{t})\ ,$ (2)
@@
$\rho_{ULDM}(r)\simeq\frac{0.019\times(m/10^{-22}\ \text{eV})^{-2}(r_{c}/\text{kpc})^{-4}}{\left[1+0.091\times(r/r_{c})^{2}\right]^{8}}\ M_{\odot}/\text{pc}^{3}\ ,$ (3)
@@
$r_{c}=0.8242\left(\frac{s}{10^{4}}\right)\left(\frac{m}{10^{-22}\ \text{eV}}\right)^{-1}\ \text{kpc}\ .$ (4)
@@
$\rho_{NFW}(r)=\frac{\rho_{s}}{\frac{r}{r_{s}}\left(1+\frac{r}{r_{s}}\right)^{2}}\ M_{\odot}/pc^{3}.$ (5)
@@
The circular velocity of a test particle moving under the influence of the spherically symmetric density profile in Eq. (2), is simply
@@
${\bf P}=\left(m,s,r_{t},r_{s},\Upsilon_{*}\right)\ .$ (7)
II.1.1 Observed Rotation Curves
[…]
At this point, as discussed in Section I, we note that we are trying to approximate a function that can take an observed rotation curve as input and infer model parameters in eq. (7) as well as their uncertainties as output. To understand how a neural network can do this, we must briefly discuss the ingredients involved in defining and training an artificial neural network in the following sub-section.
II.2 Artificial neural networks
II.2.1 Basics of neural networks
@@
$v=a({\bm{\omega}}\cdot{\bf x}+\beta)\ .$ (9)
Here, components of ${\bm{\omega}}$ are called weights and $\beta$ is called bias. The function $a(z)$ is called the activation function, which imparts a non-linearity to the transformed input. The choice of the activation function depends on the kind of problem at hand. For instance, in the case of classification problems where the required output is discreet, the sigmoid function, $a(z)=(1+e^{-z})^{-1}$ is useful since the output is contained between 00 and $1$. For regression problems, where the required output is continuous, the rectified linear unit (ReLU), $a(z)=\mathrm{max}(z,0)$ can be used.
One can also define a layer (often called a hidden layer) of $N$ neurons, where the input for each neuron is the same albeit with different weights and biases. The output of the layer can be written as a vector of size $N$, where each component is given by Eq. (9) with different weights and biases,
${\bf v}={\bf a}\left({\bm{\omega}}{\bf x}+{\bm{\beta}}\right)\ .$ (10)
In this case, ${\bm{\omega}}$ denotes a matrix of size $N\times N_{0}$ while $\beta$ is a $N\times 1$ column matrix. Also note that activation function vector ${\bf a}$ is applied element-wise.
@@
${\bf v}_{j}={\bf a}\left({\bm{\omega}_{j}}{\bf v}_{j-1}+{\bm{\beta}}_{j}\right)\ .$ (11)
Here, $j=1,2,...,L$, while ${\bm{\omega}}_{j}$ and ${\bm{\beta}}_{j}$ are the weights and biases for the $j$th layer. For ease of notation, one can denote weights and biases for all layers by $\bm{\Omega}$.
The final output of a neural network, can be written as a composition of multiple of functions ${\bf f}({\bf x},{\bf\Omega})={\bf v}_{L}({\bf v}_{L-1}(...{\bf v}_{1}({\bf x},{\bf\Omega})))$, where $L$ is the total number of layers in the neural network [38].
It must be noted that one can use $j=0$ to denote the input layer, which is just the input vector ${\bf x}$. For ANNs, no transformations occur at this layer. One can also define $j=L+1$ as the output layer, whose size will be the same as ${\bf y}$. Unlike the input layer, going from $j=L$ to $j=L+1$ does involve a transformation similar to Eq. (9), but with $a(z)=z$.
@@
In our case, the neural network of interest is shown in figure 1, where the velocities for $N_{obs}$ radius values are in the input layer, while the model parameters in eq. (7) are in the output layer, and we allow for more than $1$ hidden layers. It is worth noting, that for the case of a heteroscedastic loss function, as we shall see in Section IV.2, the size of the output layer will be doubled. This is because we shall require the uncertainties in the parameter values to be learned during training itself.
Figure 1: A schematic of the neural network that we want to construct, where given a rotation curve of dimension $M$ for a galaxy as input, one can obtain the corresponding $p$ ULDM parameters. In our work, while the number of parameters to predict will be $5$ (Eq. (7)), the number of output neurons $p=5$ or $p=10$ depending on the loss-function used.
It is also important to note that the number of observed data points in the rotation curves of different galaxies will not be the same, implying that the size of the input layer will be different for each galaxy. Further, the range of radius values for which there are observed velocities will also be different for each galaxy. Due to this, one must define a different neural network for every galaxy in our sample, i.e., we train a total of $7$ neural networks for every case in Section III and IV.
II.2.2 Training a neural network
The goal of training a neural network is to find the optimum values of the IAPs or weights and biases ${\bf\Omega}$, such that for an input from the training data, the neural network output ${\bf f}$ (also called predicted or inferred output) is close to the target output ${\bf y}$ (also called ground truth). To quantify this closeness, one employs a loss function $\mathcal{L}({\bm{\Omega}})$, which assigns a real number to each pair of predicted and target output. Training thus involves finding a set of ${\bm{\Omega}}$ such that $\mathcal{L}$ is minimized. Well-known examples of loss functions are the mean-squared error (regression tasks) and binary cross-entropy (classification tasks). In this paper, we shall consider two different loss functions: (a) mean-squared error (MSE) in Eq. (12) and (b) heteroscedastic loss in Eq. (15). The mean-squared-error loss is simply the Euclidean distance between the predictions and the target values, averaged over the number of samples, given by
$\mathcal{L}=\frac{1}{np}\sum_{i=1}^{n}|{\bf y}_{i}-{\bf\hat{y}_{i}}|^{2}\ ,$ (12)
@@ II.2.2 Training a neural network
The goal during training is to find the global minimum of the loss function in the $\bm{\Omega}$-space  [^1] This can be a high-dimensional space, since the number of IAPs range from a few thousand to tens of millions.. The usual way involves utilizing a gradient descent algorithm which requires two ingredients: (a) an efficient way to calculate the gradient of the loss function w.r.t all $\bm{\Omega}$ (i.e. weights and biases), (b) a rule to update the weights and biases in the direction of the steepest descent.
@@
We would like to point to our reader Refs.[38, 6, 40] for excellent pedagogical discussions on the various aspects of machine learning and neural networks. For our work, the choice of the gradient descent algorithm, learning rate, and other parameters related to the optimization of the loss function for our problem are discussed in Section II.5.
II.3 Generating simulated rotation curves
@@
    We make a random draw from each parameter distribution, which will give a parameter vector P (see Eq. (7)) in the $5$D parameter space.
@@
    We use $\{m,s\}$ to obtain the ULDM density profile from Eq. (3), and then use $\{r_{t},r_{s}\}$ to obtain the profile of the NFW skirt. Finally, given the total density profile, we calculate the dark matter velocity profile $V_{DM}$ using Eq. (6).
@@
    We finally construct the full velocity curve between $R_{min}$ and $R_{max}$ for all the $N_{obs}$ radius values using Eq. (1).
The ranges of the uniform distribution of each parameter are shown in Table 1 for all galaxies. We employ the above procedure for $I=5\times 10^{5}$ randomly chosen parameter vectors ${\bf P}$ from the above-mentioned ranges to obtain simulated rotation curves $v(r)\in\mathbb{R}^{N_{obs}}$. These simulated rotation curves will serve as the training inputs in Fig. 1, while ${\bf P}$ will be the target parameters that the neural networks attempts to predict for every input rotation curve.
@@
II.4 Pre-processing
Pre-processing of training data usually involves converting it to some more usable type or to normalize the input data such that all inputs have a similar range of values [44]. This is required to ensure that no component of the input is considered to be more important if it has a larger absolute value or variation. It also prevents the gradient descent algorithms from taking too small or too large steps based on the absolute value of the inputs.

### v2

@@ 1 Introduction
These novel tools and techniques are currently being applied to a wide range of astrophysical and cosmological datasets [9, 10, 11, 12, 13, 14, 15, 17, 16, 18, 19, 20, 21, 22, 23, 24]. For instance, Refs. [9, 14] use ANNs to compute likelihood in Bayesian inference to reduce computational time. On the other hand, Ref. [10] carries out non-parametric reconstruction of the Hubble parameter as a function of redshift while Ref. [19] does so for galactic rotational velocity as a function of radius. Ref. [10] finds that if one conducts Bayesian inference using Markov Chain Monte-Carlo (MCMC) on the reconstructed data, the resultant cosmological parameter posteriors are in agreement to those obtained from observed data. Deep learning has also been applied to the reconstruction of CMB B-modes[20], as well as to reconstruct full CMB spectra from partial sky data [17, 16]. Neural networks have also been used to estimate parameters from observational data like $H(z)$ data [15], CMB angular power spectrum [11], CMB birefringence maps [21] as well as Lyman-$\alpha$ spectra [22].
In this work, we use neural networks to extract model parameters from galactic rotation curves from the Spitzer Photometry & Accurate Rotation Curves (SPARC) catalog [25]. Recall that rotation curves are pivotal in tracing the mass distribution of dark matter and baryons in galaxies, and are an important test of dark matter models [26, 27]. We consider dark matter to comprise of a spin-zero particle with mass $m\sim 10^{-22}\ \text{eV}$ (called Ultra-Light Dark Matter (ULDM)), whose large deBroglie wavelength leads to the formation of a flat density core surrounded by a cold dark matter-like envelope [28] (see also [29, 30, 31]). We thus have the following five free parameters: mass of the dark matter particle m(eV)m\ ($\text${eV}), along with galaxy specific parameters such as the scaling parameter $s$, which characterizes the dark matter core, core-to-envelope transition radius $r_{t}\ (\text{kpc})$ and NFW scale radius $r_{s}\ (\text{kpc})$ which characterize the surrounding halo. The effect of Baryons is parameterized by the stellar mass-to-light ratio $\Upsilon_{*}\ (M_{\odot}/L_{\odot})$, which tunes contribution from the stellar disk (see section 2.1 for details).
@@
We explore whether, using simulated rotation curves, one can train a neural network to approximate this function. Typically, $N_{obs}\sim\mathcal{O}(15)$ for the galaxies we consider, while the number of parameters we want to infer is $5$ (or $10$ if uncertainties are also inferred). If the neural network has $2$ hidden layers with $200$ neurons each, it will have $\sim(40-50)\times 10^{3}$ internal adjustable parameters (called weights and biases). To fix these internal parameters, we need to train the neural network. For training, we use simulated rotation curves whose parameter values are already known. We generate training data for $7$ dwarf galaxies from the SPARC catalog [25] and train a different neural network for each galaxy. The size of the training data, i.e., the number of known pairs of rotation curves and parameters in our work is $\sim 10^{5}$. The details of neural networks used and their architectures are discussed in Sections 2.2.1 and 2.5 respectively.
To test our trained neural networks, we use the observed rotation curves for the $7$ galaxies as input and infer parameter values in section 3.1. Then, in section 3.2, we explore the effect of noise in the training data on the performance of the neural network during parameter inference, and find that including noise improves point-estimates of parameters when confronted with observed rotation curves, i.e., the rotation curve obtained from these parameter values agree well with the observed rotation curves. In section 4, we also utilize two different ways of obtaining uncertainties in the model parameters: during inference (as carried out in [11]), or during training (following the work in [15]). Finally, we compare the parameter point-estimates and uncertainties obtained using our approach to those obtained using MCMC in section 5. We conclude in section 6.
2 Rotation curves and artificial neural networks
2.1 Model and data
Galactic rotation curves, i.e. orbital velocity of stars and gas as a function of distance from the centres of galaxies are an important probe of the matter (visible and dark) distribution in said galaxies [26].
@@ 2.1 Model and data
$V_{obs}=\sqrt{V_{DM}^{2}+V_{g}|V_{g}|+\Upsilon_{d}V_{d}|V_{d}|+\Upsilon_{b}V_{b}|V_{b}|}\ ,$ (2.1)
@@
In this work, we consider dark matter to comprise of ultralight spin-zero scalars, with $m\sim 10^{-22}\ \text{eV}$. Due to the large deBroglie wavelength ($\lambda_{dB}\sim\mathcal{O}(\text{kpc})$), simulations suggest that dark matter halos in the ULDM paradigm have a core-halo structure where, the inner regions of the halo are described by flat density cores [28]. These cores are stationary state solutions of the Schrödinger-Poisson system of equations. In the outer regions, beyond a transition radius, ULDM behaves like CDM and the corresponding density profile can be described by the well known Navarro-Frenk-White (NFW) profile [37]. Hence, for a galactic halo, the total dark matter density profile can be written as
$\rho_{DM}(r)=\rho_{ULDM}\Theta(r_{t}-r)+\rho_{NFW}\Theta(r-r_{t})\ ,$ (2.2)
@@
$\rho_{ULDM}(r)\simeq\frac{0.019\times(m/10^{-22}\ \text{eV})^{-2}(r_{c}/\text{kpc})^{-4}}{\left[1+0.091\times(r/r_{c})^{2}\right]^{8}}\ M_{\odot}/\text{pc}^{3}\ ,$ (2.3)
@@
$r_{c}=0.8242\left(\frac{s}{10^{4}}\right)\left(\frac{m}{10^{-22}\ \text{eV}}\right)^{-1}\ \text{kpc}\ .$ (2.4)
@@
$\rho_{NFW}(r)=\frac{\rho_{s}}{\frac{r}{r_{s}}\left(1+\frac{r}{r_{s}}\right)^{2}}\ M_{\odot}/pc^{3}.$ (2.5)
@@
The circular velocity of a test particle moving under the influence of the spherically symmetric density profile in eq. (2.2), is simply
$v(r)=\sqrt{\frac{GM(r)}{r}}=\sqrt{\frac{4\pi G\int_{0}^{r}\rho_{DM}(r^{\prime})r^{\prime 2}dr^{\prime}}{r}}\ .$ (2.6)
${\bf P}=\left(m,s,r_{t},r_{s},\Upsilon_{*}\right)\ .$ (2.7)
2.1.1 Observed Rotation Curves
@@
Since ULDM affects dark matter distribution in the inner regions of galaxies, one must look at galaxies where baryons are not the dominant component even at small radii, or dark matter dominated galaxies. To study ULDM, authors in Ref. [34] chose $17$ dark matter dominated dwarf galaxies from the SPARC catalog with well-defined inner regions. In this work, we choose a subset of $7$ galaxies from the sample of $17$. It is important to note that, along with observed rotation curves, the SPARC catalog provides values for $V_{d}$ and $V_{g}$ by fitting relevant stellar density profiles. We shall utilize these values directly in eq. (2.1), while allowing $\Upsilon_{*}$ to vary.
At this point, as discussed in section 1, we note that we are trying to approximate a function that can take an observed rotation curve as input and infer model parameters in eq. (2.7) as well as their uncertainties as output. To understand how a neural network can do this, we must briefly discuss the ingredients involved in defining and training an artificial neural network in the following sub-section.
2.2 Artificial neural networks
2.2.1 Basics of neural networks
@@
$v=a({\boldsymbol{\omega}}\cdot{\bf x}+\beta)\ .$ (2.9)
Here, components of $\boldsymbol{\omega}$} are called weights and $\beta$ is called bias. The function $a(z)$ is called the activation function, which imparts a non-linearity to the transformed input. The choice of the activation function depends on the kind of problem at hand. For instance, in the case of classification problems where the required output is discreet, the sigmoid function, $a(z)=(1+e^{-z})^{-1}$ is useful since the output is contained between 0 and 11. For regression problems, where the required output is continuous, the rectified linear unit (ReLU), $a(z)=\mathrm{max}(z,0)$ can be used.
One can also define a layer (often called a hidden layer) of $N$ neurons, where the input for each neuron is the same albeit with different weights and biases. The output of the layer can be written as a vector of size $N$, where each component is given by eq. (2.9) with different weights and biases,
${\bf v}={\bf a}\left({\boldsymbol{\omega}}{\bf x}+{\boldsymbol{\beta}}\right)\ .$ (2.10)
In this case, $\boldsymbol{\omega}$} denotes a matrix of size $N\times N_{0}$ while $\beta$ is a $N\times 1$ column matrix. Also note that activation function vector $\bf a$} is applied element-wise in the above equation.
@@
${\bf v}_{j}={\bf a}\left({\boldsymbol{\omega}_{j}}{\bf v}_{j-1}+{\boldsymbol{\beta}}_{j}\right)\ .$ (2.11)
Here, j=1,2,,Lj=1,2,...,L, while ${\boldsymbol{\omega}}_{j}$ and ${\boldsymbol{\beta}}_{j}$ are the weights and biases for the $j$th layer. For ease of notation, one can denote weights and biases for all layers by $\boldsymbol{\Omega}$.
The final output of a neural network, can be written as a composition of multiple functions ${\bf f}({\bf x},{\bf\Omega})={\bf v}_{L}({\bf v}_{L-1}(...{\bf v}_{1}({\bf x},{\bf\Omega})))$, where $L$ is the total number of layers in the neural network [44].
It must be noted that one can use $j=0$ to denote the input layer, which is just the input vector $\bf x$}. For ANNs, no transformations occur at this layer. One can also define $j=L+1$ as the output layer, whose size will be the same as $\bf y$}. Unlike the input layer, going from $j=L$ to $j=L+1$ does involve a transformation similar to eq. (2.9), but with $a(z)=z$.
@@
In our case, the neural network of interest is shown in figure 1, where the velocities for $N_{obs}$ radius values are in the input layer, while the model parameters in eq. (2.7) are in the output layer, and we allow for more than $1$ hidden layers. It is worth noting, that for the case of a heteroscedastic loss function, as we shall see in section 4.2, the size of the output layer will be doubled. This is because we shall require the uncertainties in the parameter values to be learned during training itself.
Figure 1: A schematic of the neural network that we want to construct, where given a rotation curve of dimension $M$ for a galaxy as input, one can obtain the corresponding $p$ ULDM parameters. In our work, while the number of parameters to predict will be 55 (eq. (2.7)), the number of output neurons $p=5$ or $p=10$ depending on the loss-function used.
It is also important to note that the number of observed data points in the rotation curves of different galaxies will not be the same, implying that the size of the input layer will be different for each galaxy. Further, the range of radius values for which there are observed velocities will also be different for each galaxy. Due to this, one must define a different neural network for every galaxy in our sample, i.e., we train a total of $7$ neural networks for every case in section 3 and 4.
2.2.2 Training a neural network
The goal of training a neural network is to find the optimum values of the IAPs or weights and biases $\bf\Omega$}, such that for an input from the training data, the neural network output $\bf f$} (also called predicted or inferred output) is close to the target output $\bf y$} (also called ground truth). To quantify this closeness, one employs a loss function $\mathcal{L}({\boldsymbol{\Omega}})$, which assigns a real number to each pair of predicted and target output. Training thus involves finding a set of $\boldsymbol{\Omega}$} such that $\mathcal{L}$ is minimized. Well-known examples of loss functions are the mean-squared error (regression tasks) and binary cross-entropy (classification tasks). In this paper, we shall consider two different loss functions: (a) mean-squared error (MSE) in eq. (2.12) and (b) heteroscedastic loss in eq. (4.1). The mean-squared-error loss is simply the proportional to the Euclidean distance between the predictions and the target values, averaged over the number of samples, given by
$\mathcal{L}=\frac{1}{np}\sum_{i=1}^{n}|{\bf y}_{i}-{\bf\hat{y}_{i}}|^{2}\ ,$ (2.12)
@@ 2.2.2 Training a neural network
The goal during training is to find the global minimum of the loss function in the $\boldsymbol{\Omega}$-space  [^1] This can be a high-dimensional space, since the number of IAPs range from a few thousand to tens of millions.. The usual way involves utilizing a gradient descent algorithm which requires two ingredients: (a) an efficient way to calculate the gradient of the loss function w.r.t $\boldsymbol{\Omega}$, (b) a rule to update the weights and biases in the direction of the steepest descent.
@@
We would like to point to our reader Refs.[44, 6, 46] for excellent pedagogical discussions on the various aspects of machine learning and neural networks. For our work, the choice of the gradient descent algorithm, learning rate, and other parameters related to the optimization of the loss function for our problem are discussed in section 2.5.
2.3 Generating simulated rotation curves
@@
    We make a random draw from each parameter distribution, which will give a parameter vector P (see eq. (2.7)) in the 55D parameter space.
@@
    We use $\{m,s\}$ to obtain the ULDM density profile from eq. (2.3), and then use $\{r_{t},r_{s}\}$ to obtain the profile of the NFW skirt. Finally, given the total density profile, we calculate the dark matter velocity profile $V_{DM}$ using eq. (2.6).
@@
    We finally construct the full velocity curve between $R_{min}$ and $R_{max}$ for all the $N_{obs}$ radius values using eq. (2.1).
The ranges of the uniform distribution for each parameter are shown in Table 1 for all galaxies. We employ the above procedure for $I=5\times 10^{5}$ randomly chosen parameter vectors $\bf P$} from the above-mentioned ranges to obtain simulated rotation curves $v(r)\in\mathbb{R}^{N_{obs}}$. These simulated rotation curves will serve as the training inputs in figure 1, while $\bf P$} will be the target parameters that the neural networks attempts to predict for every input rotation curve.
@@
2.4 Pre-processing
Pre-processing of training data usually involves converting it to a more usable type or to normalize the input data such that all inputs have a similar range of values [50]. This is required to ensure that no component of the input is considered to be more important if it has a larger absolute value or variation. It also prevents the gradient descent algorithms from taking too small or too large steps based on the absolute value of the inputs.